OpenStudy (anonymous):

can anyone help me to integrate this respect to x; (x-1)^3. i've forgotten how to do it.

hartnn (hartnn):

u know this basic integration formula : \(\huge \int x^ndx=?\)

OpenStudy (anonymous):

i know that, but this one has power. do we separate the x and 1 and integrate them separately?

OpenStudy (anonymous):

it will follow the same formula as you would do for x^3

OpenStudy (anonymous):

3(x-1)^2 ?

OpenStudy (anonymous):

@anybody integration not differentiation

OpenStudy (turingtest):

will somebody explain why we need a u-sub please, I think that is where the asker is having difficulty

OpenStudy (anonymous):

\[\frac{ (x-1)^{3} }{ 4 }\]?

OpenStudy (turingtest):

\[\int(x-1)^3dx\]now notice that dx is the differential of the argument x-1, so we call it u: \[u=x-1\implies du=dx\]after the sub the integral is then\[\int u^3du\]can you integrate that?

OpenStudy (anonymous):

u^4/4

OpenStudy (turingtest):

great, now sub back in for what u is...

OpenStudy (turingtest):

oh and always include the +C if it is indefinite

OpenStudy (anonymous):

where do we get the +c?

OpenStudy (turingtest):

that's a different question, all indefinite integrals (integrals where the bounds are not specified) need a +C at the end where C is some unknown constant if you want an explanation for that post it separately please for now just sub in what we had fopr u into u^4/4 and add a +C at the end and tell me what you get

OpenStudy (anonymous):

(x-1)^4/4

OpenStudy (turingtest):

+C ;)

OpenStudy (anonymous):

okey, thanks ;)

OpenStudy (turingtest):

welcome if you want me to explain the +C part I can try, but it is a little subtle, and please post it as a separate question welcome!